If the inventor in exercise 53(see below) charges $4 per unit,then her profit for producing and selling x units is given by the function
P(x) =2.25x-7000
a. what is her profit if she sells 2000 units?
b. Waht is her profti if she sells 5000 units?
c. what is the break-even point for sales?
53. from above
Th inventor of a new product believes that the cost of producing the product is given by funciton C(x)=1.75x+7000
Mcgraw hill 6th edition, chapter 7.5, exercise # 54.
Thanks fo much for your valuable time!!!

Answer by Hook(36)

Don't worry about what problem number 53 says. Here's what you've got to do:
a. what is her profit if she sells 2000 units?
-Find P(2000). Plug 2000 in for x in your profit function and see what you get:
P(2000) =2.25(2000)-7000
b. what is her profit if she sells 5000 units?
-Nearly the same as a. Find P(5000). Plug 5000 in for x in your profit function and see what you get:
P(5000) =2.25(5000)-7000
c. what is the break-even point for sales?
-Set P(x) equal to 0 and solve for x. Like this:
2.25x - 7000 = 0